OpenStudy (princesssleelee):

Let f(x) = x2 − 2x + 2 and g(x) = x − 3. Find f(x) ⋅ g(x). x3 − 6x2 + 12x − 6 x3 − 5x2 + 8x − 6 x3 − 3x2 + 6x − 6 x3 − 2x2 + 5x − 6

OpenStudy (princesssleelee):

I am thinking Letter D...

OpenStudy (anonymous):

\[=f(x)\times~g(x)\]\[=(x^2-2x+2)(x-3)\]

OpenStudy (princesssleelee):

Can you tell me how you configured that? please.

OpenStudy (olivia567):

Just multiply 2 equations.

OpenStudy (anonymous):

\[=(x^2-2x+2)(x-3)\]what do u get when u multiply x^2 with x?

OpenStudy (princesssleelee):

x^3

OpenStudy (anonymous):

yes :) so what is \[x^2\times-3=?\]

OpenStudy (princesssleelee):

-3x^2

OpenStudy (princesssleelee):

making the answer c, correct? :)

OpenStudy (anonymous):

Let me check

OpenStudy (anonymous):

nope,C is not the answer\[=(x^2-2x+2)(x-3)\]\[=x^3-3x^2-2x^2+6x+2x-6\]\[=x^3-3x^2+8x-6\]

OpenStudy (princesssleelee):

B! I figured it out as well lol

OpenStudy (princesssleelee):

Did not know to use FOIL!

OpenStudy (princesssleelee):

Did you get letter choice B as well?

OpenStudy (anonymous):

yes,the answer is B :)

OpenStudy (princesssleelee):

Thank you both!

OpenStudy (princesssleelee):

Could you check one more of my questions, maybe two? Lol

OpenStudy (anonymous):

sure

OpenStudy (princesssleelee):

Let f(x) = x + 7 and g(x) = x − 4. Find f(x) ⋅ g(x). x2 − 3x − 28 x2 − 3x − 11 x2 + 3x − 28 x2 + 3x − 11

OpenStudy (princesssleelee):

I got C for this one

OpenStudy (olivia567):

Welcome.. :) What are the next questions?

OpenStudy (anonymous):

correct!

OpenStudy (princesssleelee):

Multiply (3x − 5)2. 9x2 − 30x + 25 6x2 − 10 9x2 − 25 9x2 − 15x + 25

OpenStudy (princesssleelee):

I got a little confused, but I got B

OpenStudy (anonymous):

\[=(3x-5)^2\]\[=(3x-5)(3x-5)\]

OpenStudy (princesssleelee):

Im thinking maybe A as well

OpenStudy (princesssleelee):

Yes, A it looks like lol

OpenStudy (olivia567):

you are right.. (y) :)

OpenStudy (princesssleelee):

Nope. D lmao

OpenStudy (princesssleelee):

Wait, A?

OpenStudy (princesssleelee):

If I foil the equation @MARC_ gave me, I get D

OpenStudy (anonymous):

\[=(3x-5)(3x-5)\]\[=9x^2-15x-15x+25\]\[=?\]

OpenStudy (olivia567):

yes.. A is right one.\[(a-b)^{2}= a ^{2}-2ab+b ^{2}\]

OpenStudy (princesssleelee):

Forgot to combine, which I did correctly the second time lol, Yes A! I confuse myself WAYY too much!

OpenStudy (princesssleelee):

Thank you both so so so much! I got a 100% Im glad you both were able to help me!! As well as I am proud of myself for getting the others correct on my own! :)

OpenStudy (princesssleelee):

Someone please give @Olivia567 a medal, she deserves it!